OpenStudy (anonymous):

Find the slope of a line that passes through (–2, –3) and (1, 1).

OpenStudy (anonymous):

@AriPotta can you help me plz??.. this is for Algebra 1

OpenStudy (aripotta):

use the slope formula

OpenStudy (anonymous):

Oh yea so it would be \[\frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

For the equation –4y = 8x, what is the constant of variation? (1 point) –4 –2 1 2

OpenStudy (anonymous):

i think it might be -2

OpenStudy (aripotta):

yep

OpenStudy (anonymous):

ok so What is an equation for the line with slope \[\frac{ 2 }{ 3 }\] and y-intercept 9

OpenStudy (aripotta):

well, you know slope-intercept form, right? just plug those into the equation

OpenStudy (anonymous):

i dont kno tht

OpenStudy (aripotta):

slope-intercept form is y = mx + b. m is the slope and b is the y-intercept

OpenStudy (aripotta):

so your answer would be y = 2/3x + 9

OpenStudy (anonymous):

Yea. I kinda guessed tht was the answer but i wasnt sure...

OpenStudy (anonymous):

What is an equation in slope-intercept form for the line that passes through the points (1, –3) and (3, 1)?

OpenStudy (aripotta):

find the slope, put it into point-slope form, then change it to slope-intercept

OpenStudy (anonymous):

im confused!

OpenStudy (aripotta):

point-slope form is y - y1 = m(x - x1)

OpenStudy (anonymous):

so would it be y = 3x + 1

OpenStudy (aripotta):

no

OpenStudy (aripotta):

3 isn't the slope

OpenStudy (anonymous):

y = 2x + 5

OpenStudy (aripotta):

y = 2x - 5

OpenStudy (anonymous):

Yay! thank you for your help! i got an A+....

OpenStudy (aripotta):

great! :) hopefully you understand it better now

OpenStudy (anonymous):

yes i do! thank you!